Two water tanks at PPV Aman and PPV Damai consists of x and 3000 litres of water, respectively, on the first day.


a) Subsequently, y litres of water is used from the tank in PPV Aman everyday. Volume of water left at the end of 18th day is 2520 litres and the tank will be expected empty by the day 81. Find the value of x and y. [8 marks]


b) The usage of water in PPV Damai constantly 30 litres per day. Find the day for both, tank at PPV Aman and PPV Damai, has the same amount of water. [4 marks]


1
Expert's answer
2021-11-09T16:58:23-0500

a)

"x-18y=2520"

"x=81y"

Then


"81y-18y=2520"

"63y=2520"

"y=40"

"x=81(40)=3240"

"x=3240, y=40"


b)


"3240-40t=3000-30t"

"10t=240"

"t=24"

Tank at PPV Aman and PPV Damai, has the same amount of water at the end of 24th day.